# Super-extensions of tensor algebras and their applications

### v1

@@ 3. Properties of operators
$L(e_{i})L(e_{j}^{\ast})(v_{d}\otimes\cdots\otimes v_{1}\otimes\sigma)$
@@
$\mathcal{S}_{1}$ =$=\{1,s_{1}\},$
$\mathcal{S}_{2}$ =$=\{1,s_{2},s_{2}s_{1}\},$
$\mathcal{S}_{3}$ =$=\{1,s_{3},s_{3}s_{2},s_{3}s_{2}s_{1}\}$
@@
$\mathcal{S}_{d-1}$ =$=\{1,s_{d-1},s_{d-1}s_{d-2},\ldots,s_{d-1}s_{d-2}\cdots s_{2}s_{1}\}.$
@@
$=\mathbf{m}(I)!e_{i_{d}}\cdots e_{i_{1}}.$
@@
$L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(e_{j_{d}}\cdots e_{j_{1}})=\begin{cases}\mathbf{m}(I)!e_{i_{d}}\cdots e_{i_{1}},&\text{if }\mathbf{m}(I)=\mathbf{m}(J),\\[3.0pt] 0,&\text{if }\mathbf{m}(I)\neq\mathbf{m}(J).\end{cases}$
@@ 4. Generalizations of the Schur-Sergeev duality
Let $W$ be a vector space and $A$ a subset of $\mathrm{End}_{\mathbb{C}}(W)$. The commutant of $A$ is denoted by $\mathrm{Comm}(A)$. That is
$\mathrm{Comm}(A)=\{f\in\mathrm{End}_{\mathbb{C}}(W)|fg=gf\text{ for all }g\in A\}.$
@@
with $w_{1},\ldots,w_{d}\in V$, $\sigma\in\mathfrak{S}_{d+k}$, and $v_{1}^{\ast},\ldots,v_{k}^{\ast}\in V^{\ast}$. For any $t\in\mathbb{C}\mathfrak{S}_{d+k}$, we define $\pi_{d+k}(t)\in\mathrm{End}_{\mathbb{C}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$ by $\pi_{d+k}(t)(\varphi)=\varphi\cdot t$ where $\varphi\in V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k}$. Denote $\mathcal{B}_{d+k}=\pi_{d+k}(\mathbb{C}\mathfrak{S}_{d+k})$.
@@
Keep the notations above. We have
$\mathrm{Comm}(\mathcal{A}_{d}^{k})=\mathcal{B}_{d+k},$
@@
$\mathrm{Comm}(\mathcal{B}_{d+k})=\mathcal{A}_{d}^{k}.$
@@
Apparently, by Proposition 2.1, we have $\mathcal{A}_{d}^{k}\subset\mathrm{Comm}(\mathcal{B}_{d+k})$. Now for any $f\in\mathrm{Comm}(\mathcal{B}_{d+k})$ and $\varphi\in V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k}$, it follows from Proposition 3.6 that
$f(\varphi)$
$=f(A_{d}(\varphi))$
@@
Hence $f\in\mathcal{A}_{d}^{k}$. As a consequence, $\mathrm{Comm}(\mathcal{B}_{d+k})\subset\mathcal{A}_{d}^{k}$.
Since $\mathbb{C}\mathfrak{S}_{d+k}$ is a semisimple algebra, it follows from the double commutant theorem that $\mathrm{Comm}(\mathcal{A}_{d}^{k})=\mathcal{B}_{d+k}$.∎
As an application of the above theorem, we provide a new proof of the Schur-Sergeev duality theorem. Firstly, we need the following technical lemma.
@@
Theorem 4.3 (Schur-Sergeev duality).
@@
$\mathrm{Comm}\big{(}\rho_{d}(U(gl(m|n)))\big{)}=\mathcal{B}_{d},$
@@
$\mathrm{Comm}(\mathcal{B}_{d})=\rho_{d}(U(gl(m|n))).$
@@
We are going to show that $\mathcal{A}_{d}^{0}=\rho_{d}(U(gl(m|n)))$. We first notice that every element in $\mathcal{A}_{d}^{0}$ is a linear combination of operators of the form
@@
Therefore $\rho_{d}(E_{ij})\in\mathcal{A}_{d}^{0}$, and hence $\rho_{d}(U(gl(m|n)))\subset\mathcal{A}_{d}^{0}$.
On the other hand, for any $(i_{p},\ldots,i_{1}),(j_{p},\ldots,j_{1})\in[m+n]^{p}$, we use induction on $p$ to show that $L(e_{i_{p}})\cdots L(e_{i_{1}})L(e_{j_{1}}^{\ast})\cdots L(e_{j_{p}}^{\ast})\in\rho_{d}(U(gl(m|n)))$. The case $p=1$ follows from (3.1). Assume that the result holds for $p\geq 1$. By the above lemma, it is easy to see that
@@ 5.1. Hecke algebras
Let $\mathscr{H}_{\infty}$ be the inductive limit of the natural inclusions $\mathscr{H}_{0}\subset\mathscr{H}_{1}\subset\cdots$. Given $k\geq 1$, we define an algebra homomorphism :${}^{\uparrow k}:\mathscr{H}_{\infty}\rightarrow\mathscr{H}_{\infty}$ by $T_{i}^{\uparrow k}=T_{i+k}$.
@@
$\gamma(i,j)=\left\{\begin{array}[]{ll}1,&\mbox{if $i>j$},\\ -1,&\mbox{if $i\leq j$}.\end{array}\right.$
@@
It can be verified by the braid relation $T_{1}T_{2}T_{1}=T_{2}T_{1}T_{2}$ case by case. ∎
@@ 5.2. $q$-tensor superspaces
$\widehat{i}=\left\{\begin{array}[]{ll}0,&\mbox{if $1\leq i\leq m$},\\ 1,&\mbox{if $m<i\leq m+n$}\end{array}\right.$
@@
Let $\mathbb{V}=\mathbb{V}_{\overline{0}}\oplus\mathbb{V}_{\overline{1}}$ be a vector space over $\mathbb{C}(q)$ with a specified basis $\{e_{1},e_{2},\ldots,e_{m+n}\}$, where $\mathbb{V}_{\overline{0}}$ (resp. $\mathbb{V}_{\overline{1}}$) is spanned by $\{e_{i}|1\leq i\leq m\}$ (resp. $\{e_{i}|m+1\leq i\leq m+n\}$).
@@
As before, $\mathfrak{S}_{d}$ acts on $[m+n]^{d}$ from the right hand side. The right $\mathscr{H}_{d}$-action on $\mathbb{V}^{\otimes d}$ is defined on the basis element $e_{I}$ as follows: for $1\leq k<d$,
(5.1) $e_{I}.T_{k}=\left\{\begin{array}[]{ll}(-1)^{\widehat{i_{k}}\widehat{i_{k+1}}}e_{I.s_{k}},&\mbox{if $i_{k}>i_{k+1}$},\\ (-1)^{\widehat{i_{k}}}q_{i_{k}}e_{I},&\mbox{if $i_{k}=i_{k+1}$},\\ (-1)^{\widehat{i_{k}}\widehat{i_{k+1}}}e_{I.s_{k}}+(q-q^{-1})e_{I},&\mbox{if $i_{k}<i_{k+1}$}.\end{array}\right.$
@@ 5.3. Super-extensions
Endow an algebra structure on $\widehat{T}(\mathbb{V})$ whose product $\cdot$ is defined by
@@ 5.4. Multiplications and Derivations
It is easy to see that $L(\varphi)$ is well-defined. For $L(e_{j}^{*})$, we only need to show that it commutes with the action of $\mathscr{H}_{\infty}$. We introduce a linear map $h_{k}^{(j)}:T_{d}(\mathbb{V})\rightarrow\widetilde{T}_{d}(\mathbb{V})$ for $1\leq k\leq d$ defined by
@@
It is just the $k$-th term of $L(e_{j}^{*})$ over $T_{d}(\mathbb{V})$ in (5.3). For any $T\in\mathscr{H}_{\infty}$, we define $\pi(T)\in\mathrm{End}_{\mathbb{C}}(\widetilde{T}(\mathbb{V}))$ by $\pi(T)(\varphi)=\varphi\cdot T$ for $\varphi\in\widetilde{T}(\mathbb{V})$. To prove the well-definedness of $L(e_{j}^{*})$, it suffices to show that $\pi(T_{1}),...,\pi(T_{d-1})$ commutes with $\sum_{k=1}^{d}h_{k}^{(j)}$. We have the following lemma:
@@
Write $h_{k}$ for $h_{k}^{(j)}$. Then for $r=1,...,d-1$, we have
@@
When $k+2\leq r\leq d-1$, we show that $\pi(T_{r})$ commutes with $h_{k}$ by induction on $k$. We start the induction by $\pi(T_{r})h_{r-1}=h_{r-1}\pi(T_{r})$ which is the result of the last paragraph. For $k<r-1$, we have
@@
Adding the above two equations, we get the result. ∎
@@ 5.5. General linear quantum supergroups
The general linear quantum supergroup $\mathbf{U}_{m|n}=U_{q}(\mathfrak{gl}_{m|n})$ is a $\mathbb{C}(q)$-algebra generated by
$\left\{\begin{array}[]{lll}\mbox{even generators:}&E_{i},F_{i},K_{j}^{\pm 1},&1\leq i,j\leq m+n,i\neq m,m+n;\\ \mbox{odd generators:}&E_{m},F_{m}&\end{array}\right.$
@@
There is a Hopf superalgebra structure on $\mathbf{U}_{m|n}$, whose comultiplication $\Delta$ is
@@
Particularly, $|\Delta^{(d)}(E_{i})|=|E_{i}|$ and $|\Delta^{(d)}(F_{i})|=|F_{i}|$, where ||$\cdot$|| ⋅ | means the parity of elements in $\mathbf{U}_{m|n}$.
The superspace $\mathbb{V}$ admits a natural representation of $\mathbf{U}_{m|n}$ via
@@
Then the $q$-tensor superspace $\mathbb{V}^{\otimes d}$ is also equipped with a $\mathbf{U}_{m|n}$-module structure via $\Delta^{(d)}$. We shall show that the $\mathbf{U}_{m|n}$-module structure can be extended over the $\widetilde{T}(\mathbb{V})$ as $x.(e_{I}\otimes T_{\sigma})=(x.e_{I})\otimes T_{\sigma}$ for $x\in\mathbf{U}_{m|n},e_{I}\in\mathbb{V}^{\otimes d},T_{\sigma}\in\mathscr{H}_{\infty}$. It can be seen instantly for $K$, since
@@ 6.1. Commutation relations
$K_{j}L(e_{i})=q_{j}^{\delta_{ij}}L(e_{i})K_{j},\quad K_{j}L(e_{i}^{*})=q_{j}^{-\delta_{ij}}L(e_{i}^{*})K_{j},\quad K_{j}L(T_{k})=L(T_{k})K_{j}$
@@
[…]
@@
6.2. The representations of $U_{m|n}$ on $\widehat{T}(\mathbb{V})$
@@
For $I=(i_{d-1},\ldots,i_{2},i_{1})\in[m+n]^{d-1},k\in[m+n]$, by direct calculation, we have
@@
We only prove the first equation. The second one can be proved similarly. We use induction on $d$. It is clear for $d=1$. Suppose the equation holds for $d=r-1$, then
@@
$=A_{i}^{(r)}(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})=L(e_{i})L(e_{i+1}^{*})(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}}),$
where the forth equality follows from the inductive hypothesis, and the sign comes from the super-action of $E_{i}K_{i}^{-1}$.
We further denote $E_{i,i+1}=E_{i}$, $E_{i+1,i}=F_{i}$ and
@@
As above, we only prove the first equation. We use induction on $j-i$. It has been showed in Lemma 6.5 for the case $j-i=1$. On $\mathbb{V}^{\otimes d}$, suppose $E_{ik}K_{i}^{-1}=L(e_{i})L(e_{k}^{*}),i<k<j$. Then
@@
Based on these identities on $\mathbb{V}^{\otimes d}$ and the fact that $L(e),L(e^{*}),K,K^{-1}$ all commute with $\mathscr{H}_{\infty}$, we can actually have a well-defined $\mathbf{U}_{m|n}$-module structure on $\widetilde{T}(\mathbb{V})$ and get the same result.
@@
$L(v_{k})\cdots L(v_{1})L(v_{1}^{*})\cdots L(v_{k}^{*})\in\rho(\mathbf{U}_{m|n}),$
where $\rho$ denotes the representation of $U_{m|n}$ on $\widetilde{T}(\mathbb{V})$.
@@
where $a,b,c,d\in\mathbb{C}(q)$. Recall from Proposition 6.1 and 6.3 that
@@
where $c_{1},c_{2}$ is ±$\pm(q-q^{-1})$ or 00. By direct calculation, we have $c_{1}=0$ or $c_{2}=0$ (i.e., $c_{1}c_{2}=0$), since $\gamma(i,j)=-\gamma(k,i)$ or $\gamma(j,i)=-\gamma(k,j)$. Then we get the claim. From the claim, we can see that

### v2

@@ 3. Properties of operators
$L(e_{i})L(e_{j}^{\ast})(v_{d}\cdots v_{1}\otimes\sigma)$
@@
$\mathcal{S}_{1}$ $=\{1,s_{1}\},$
$\mathcal{S}_{2}$ $=\{1,s_{2},s_{2}s_{1}\},$
$\mathcal{S}_{3}$ $=\{1,s_{3},s_{3}s_{2},s_{3}s_{2}s_{1}\},$
@@
$\mathcal{S}_{d-1}$ $=\{1,s_{d-1},s_{d-1}s_{d-2},\ldots,s_{d-1}s_{d-2}\cdots s_{2}s_{1}\}.$
@@
$=\mathbf{m}(I)!e_{i_{d}}\cdots e_{i_{1}},$
where the fifth equality follows by an isomorphism $(\mathfrak{S}_{d})_{I.\tau}\to(\mathfrak{S}_{d})_{I},\sigma\mapsto\tau\sigma\tau^{-1}$. ∎
@@
$L(e_{i_{d}})\cdots L(e_{i_{1}})L(e_{i_{1}}^{\ast})\cdots L(e_{i_{d}}^{\ast})(e_{j_{d}}\cdots e_{j_{1}})=\begin{cases}\mathbf{m}(I)!e_{i_{d}}\cdots e_{i_{1}}&\text{if }\mathbf{m}(I)=\mathbf{m}(J),\\[3.0pt] 0&\text{if }\mathbf{m}(I)\neq\mathbf{m}(J).\end{cases}$
@@ 4. Generalizations of the Schur-Sergeev duality
Let $W$ be a vector space and $A$ a subset of $\mathrm{End}_{\mathbb{C}}(W)$. The commutant of $A$ is denoted by $\mathrm{End}_{A}(W)$. That is
$\mathrm{End}_{A}(W)=\{f\in\mathrm{End}_{\mathbb{C}}(W)~|~fg=gf\text{ for all }g\in A\}.$
@@
with $w_{1},\ldots,w_{d}\in V$, $\sigma\in\mathfrak{S}_{d+k}$, and $v_{1}^{\ast},\ldots,v_{k}^{\ast}\in V^{\ast}$. The representation $\pi_{d+k}$ of $\mathbb{C}\mathfrak{S}_{d+k}$ on $V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k}$ is given by right multiplication and we identify $\pi_{d+k}(\mathbb{C}\mathfrak{S}_{d+k})$ with $\mathbb{C}\mathfrak{S}_{d+k}$.
@@
We have
$\mathcal{A}_{d}^{k}=\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d+k}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$
@@
$\mathrm{End}_{\mathcal{A}_{d}^{k}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})=\mathbb{C}\mathfrak{S}_{d+k}.$
@@
Apparently, by Proposition 2.1, we have $\mathcal{A}_{d}^{k}\subset\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d+k}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$. Now for any $f\in\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d+k}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})$ and $\varphi\in V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k}$, it follows from Proposition 3.6 that
$f(\varphi)$ $=f(A_{d}(\varphi))$
@@
Hence $f\in\mathcal{A}_{d}^{k}$. As a consequence, $\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d+k}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})\subset\mathcal{A}_{d}^{k}$.
Since $\mathbb{C}\mathfrak{S}_{d+k}$ is a semisimple algebra, it follows from the double commutant theorem that $\mathrm{End}_{\mathcal{A}_{d}^{k}}(V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d+k})=\mathbb{C}\mathfrak{S}_{d+k}$.∎
As an application of the above theorem, we provide a new proof of the Schur-Sergeev duality. Firstly, we need the following technical lemma.
@@
Corollary 4.3 (Schur-Sergeev duality).
@@
$\rho_{d}(U(\mathfrak{gl}(m|n)))=\mathrm{End}_{\mathbb{C}\mathfrak{S}_{d}}(V^{\otimes d})$
@@
$\mathrm{End}_{\rho_{d}(U(\mathfrak{gl}(m|n)))}(V^{\otimes d})=\mathbb{C}\mathfrak{S}_{d},$
where $\rho_{d}$ denotes the representation of $U(\mathfrak{gl}(m|n))$ on $V^{\otimes d}\cong V^{\otimes d}\otimes_{\mathbb{C}\mathfrak{S}_{d}}\mathbb{C}\mathfrak{S}_{d}$.
@@
We are going to show that $\mathcal{A}_{d}^{0}=\rho_{d}(U(\mathfrak{gl}(m|n)))$. We first notice that every element in $\mathcal{A}_{d}^{0}$ is a linear combination of operators of the form
@@
Therefore $\rho_{d}(E_{ij})\in\mathcal{A}_{d}^{0}$, and hence $\rho_{d}(U(\mathfrak{gl}(m|n)))\subset\mathcal{A}_{d}^{0}$.
On the other hand, for any $(i_{p},\ldots,i_{1}),(j_{p},\ldots,j_{1})\in[m+n]^{p}$, we use induction on $p$ to show that $L(e_{i_{p}})\cdots L(e_{i_{1}})L(e_{j_{1}}^{\ast})\cdots L(e_{j_{p}}^{\ast})\in\rho_{d}(U(\mathfrak{gl}(m|n)))$. The case $p=1$ follows from (3.1). Assume that the result holds for $p\geq 1$. By the above lemma, it is easy to see that
@@ 5.1. Hecke algebras
Let $\mathscr{H}_{\infty}$ be the inductive limit of the natural inclusions $\mathscr{H}_{0}\subset\mathscr{H}_{1}\subset\cdots$. Given $k\geq 1$, we define an algebra homomorphism $^{\uparrow k}$$\mathscr{H}_{\infty}\rightarrow\mathscr{H}_{\infty}$ by $T_{i}^{\uparrow k}=T_{i+k}$.
@@
$\gamma(i,j)=\left\{\begin{array}[]{ll}1&\mbox{if $i>j$},\\ -1&\mbox{if $i\leq j$}.\end{array}\right$.
@@
It can be verified directly by the braid relation $T_{1}T_{2}T_{1}=T_{2}T_{1}T_{2}$ case by case. ∎
@@ 5.2. $q$-tensor superspaces
$\widehat{i}=\left\{\begin{array}[]{ll}0&\mbox{if $1\leq i\leq m$},\\ 1&\mbox{if $m<i\leq m+n$}\end{array}\right$.
@@
Let $\mathbb{V}=\mathbb{V}_{\overline{0}}\oplus\mathbb{V}_{\overline{1}}$ be a vector space over $\mathbb{C}(q)$ with a specified basis $\{e_{1},e_{2},\ldots,e_{m+n}\}$, where $\mathbb{V}_{\overline{0}}$ (resp. $\mathbb{V}_{\overline{1}}$) is spanned by $\{e_{i}~|~1\leq i\leq m\}$ (resp. $\{e_{i}~|~m+1\leq i\leq m+n\}$).
@@
As before, $\mathfrak{S}_{d}$ acts on $m+n]^{d}$ from the right hand side. The right $\mathscr{H}_{d}$-action on $\mathbb{V}^{\otimes d}$ is defined on the basis element $e_{I}$ as follows: for $1\leq k<d$,
(5.1) $e_{I}.T_{k}=\left\{\begin{array}[]{ll}(-1)^{\widehat{i_{k}}\widehat{i_{k+1}}}e_{I.s_{k}}&\mbox{if $i_{k}>i_{k+1}$},\\ (-1)^{\widehat{i_{k}}}q_{i_{k}}e_{I}&\mbox{if $i_{k}=i_{k+1}$},\\ (-1)^{\widehat{i_{k}}\widehat{i_{k+1}}}e_{I.s_{k}}+(q-q^{-1})e_{I}&\mbox{if $i_{k}<i_{k+1}$}.\end{array}\right$.
@@ 5.3. Super-extensions
Endow an algebra structure on $\widetilde{T}(\mathbb{V})$ whose product $\cdot$ is defined by
@@ 5.4. Multiplications and Derivations
It is easy to see that $L(\varphi)$ is well-defined. For $L(e_{j}^{*})$, we only need to show that it commutes with the action of $\mathscr{H}_{\infty}$. We introduce a linear map $h_{k}^{(j)}:\widetilde{T}_{d}(\mathbb{V})\rightarrow\widetilde{T}_{d}(\mathbb{V})$ for $1\leq k\leq d$ defined by
@@
It is just the $k$-th term of $L(e_{j}^{*})$ on $\widetilde{T}_{d}(\mathbb{V})$ in (5.3). For any $T\in\mathscr{H}_{\infty}$, we define $\pi(T)\in\mathrm{End}_{\mathbb{C}}(\widetilde{T}(\mathbb{V}))$ by $\pi(T)(\varphi)=\varphi\cdot T$ for $\varphi\in\widetilde{T}(\mathbb{V})$. To prove the well-definedness of $L(e_{j}^{*})$, it suffices to show that $\pi(T_{1}),...,\pi(T_{d-1})$ commutes with $\sum_{k=1}^{d}h_{k}^{(j)}$.
@@
Write $h_{k}$ for $h_{k}^{(j)}$. Then for r=1,,d1r=1,...,d-1, we have
@@
When $k+2\leq r\leq d-1$, we show that $\pi(T_{r})$ commutes with $h_{k}$ by induction on $k$. We start the induction by $\pi(T_{r})h_{r-1}=h_{r-1}\pi(T_{r})$ which is the result of the last paragraph. For k<r1k<r-1, we have
@@
Adding up the above two equations, we get the result. ∎
@@ 5.5. General linear quantum supergroups
The general linear quantum supergroup $U_{q}(\mathfrak{gl}(m|n))$ is a $\mathbb{C}(q)$-algebra generated by
[…]
@@
There is a Hopf superalgebra structure on $U_{q}(\mathfrak{gl}(m|n))$, whose comultiplication $\Delta$ is
@@
Particularly, $|\Delta^{(d)}(E_{i})|=|E_{i}|$ and $|\Delta^{(d)}(F_{i})|=|F_{i}|$, where $|\cdot|$ means the parity of elements in $U_{q}(\mathfrak{gl}(m|n))$.
The superspace $\mathbb{V}$ admits a natural representation of $U_{q}(\mathfrak{gl}(m|n))$ via
@@
Then the $q$-tensor superspace $\mathbb{V}^{\otimes d}$ is also equipped with a $U_{q}(\mathfrak{gl}(m|n))$-module structure via $\Delta^{(d)}$. We shall show that the $U_{q}(\mathfrak{gl}(m|n))$-module structure can be extended to $\widetilde{T}(\mathbb{V})$ by $x.(e_{I}\otimes T_{\sigma})=(x.e_{I})\otimes T_{\sigma}$ for any $x\in U_{q}(\mathfrak{gl}(m|n)),e_{I}\in\mathbb{V}^{\otimes d},T_{\sigma}\in\mathscr{H}_{\infty}$. It can be seen instantly for $K$, since
@@ 6.1. Commutation relations
$K_{j}L(e_{i})=q_{j}^{\delta_{ij}}L(e_{i})K_{j},\quad K_{j}L(e_{i}^{*})=q_{j}^{-\delta_{ij}}L(e_{i}^{*})K_{j},\quad K_{j}L(T_{k})=L(T_{k})K_{j}.$
@@
(6.4) $L(e_{i}^{*})L(e_{j}^{*})|_{\widetilde{T}_{d}(\mathbb{V})}=(-1)^{\widehat{i}\widehat{j}}q_{i}^{\delta_{ij}}L(T_{1}^{\gamma(i,j)})L(e_{j}^{*})L(e_{i}^{*})|_{\widetilde{T}_{d}(\mathbb{V})},$
@@
6.2. The representations of $U_{q}(\mathfrak{gl}(m|n))$ on $\widetilde{T}(\mathbb{V})$
@@
For $I=(i_{d-1},\ldots,i_{2},i_{1})\in[m+n]^{d-1}$ and $k\in[m+n]$, by direct calculation, we have
@@
We only prove the first equation. The second one can be proved similarly. We use induction on $d$. It is clear for $d=1$. Suppose the equation holds for d=r1d=r-1, then
@@
$=A_{i}^{(r)}(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}})$
$=L(e_{i})L(e_{i+1}^{*})(e_{i_{r}}e_{i_{r-1}}\cdots e_{i_{1}}),$
where the forth equality follows from the inductive hypothesis, and the sign comes from the super-action of $E_{i}K_{i}^{-1}$.∎
We further denote $E_{i,i+1}=E_{i}$, $E_{i+1,i}=F_{i}$, and
@@
As above, we only prove the first equation. We use induction on jij-i. It has been showed in Lemma 6.5 for the case ji=1j-i=1. On $\mathbb{V}^{\otimes d}$, suppose $E_{ik}K_{i}^{-1}=L(e_{i})L(e_{k}^{*})$ for $i<k<j$. Then
@@
Based on these identities on $\mathbb{V}^{\otimes d}$ and the fact that $L(e),L(e^{*}),K,K^{-1}$ all commute with $\mathscr{H}_{\infty}$, we can actually have a well-defined $U_{q}(\mathfrak{gl}(m|n))$-module structure on $\widetilde{T}(\mathbb{V})$ and get the same result.
@@
$L(v_{k})\cdots L(v_{1})L(v_{1}^{*})\cdots L(v_{k}^{*})\in\rho(U_{q}(\mathfrak{gl}(m|n))),$
where $\rho$ denotes the representation of $U_{q}(\mathfrak{gl}(m|n))$ on $\widetilde{T}(\mathbb{V})$.
@@
where $a,b,c,d\in\mathbb{C}(q)$. Recall from Propositions 6.1 and 6.3 that
@@
where $c_{1},c_{2}$ is $\pm(q-q^{-1})$ or 0. By direct calculation, we have $c_{1}=0$ or $c_{2}=0$ (i.e., $c_{1}c_{2}=0$), since $\gamma(i,j)=-\gamma(k,i)$ or $\gamma(j,i)=-\gamma(k,j)$. Then we get the claim.
From the claim, we can see that